OpenStudy (anonymous):

Dalco Manufacturing estimates that its weekly profit, P, in hundreds of dollars, can be approximated by the formula P= -2x^2 + 8x + 2, where x is the number of units produced per week, in thousands. a.How many units should the company produce per week to earn the maximum profit? b.Find the maximum weekly profit

OpenStudy (anonymous):

PART A: To find the maximum point of any function we need to know where it's derivative equals zero: \[P(\max) \rightarrow \frac{ dP }{ dx }=0 \] So to start we need to find the derivative: \[\frac{ dP }{ dx } = \frac{d}{dx}(-2x^{2}+8x+2)=-4x+8\] We can then find the maximum by setting this equal to zero: \[-4x+8=0\] and solving for x \[4x=8\]\[x=2\] So, for maximum profit they should produce 2 units/week. PART B We know how man units they need to produce and we know the profit function. To find the profit amount we need to evaluate the profit function at x=2. \[-2x^{2}+8x+2\]\[-2(2)^{2}+8(2)+2\]\[10\] The maximum profit will be 10 * $100 = $1,000/week when 2 units are produced.

OpenStudy (anonymous):

Is there a simpler way to explain this to me? @twvogels

OpenStudy (anonymous):

Maybe. What part don't you understand?

OpenStudy (anonymous):

For starters: To find the maximum point of any function we need to know where it's derivative equals zero: P(max)→dP/dx=0

OpenStudy (anonymous):

Have you taken Calculus 1 yet? (limits, derivatives etc.) That's the fastest way to solve it but there are other ways to do it.

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

Hmm...okay. You can also solve it by graphing the function. If you do that you can read the maximum right off the chart.

OpenStudy (anonymous):

It will need to be fairly accurate though. Use a calculator or computer, it will be pretty hard to do it neatly by hand.

OpenStudy (anonymous):

I have graphed it, but I don't get an answer that makes since. I keep getting a negative.

OpenStudy (anonymous):

Okay hold on. I'll graph it and see what I get. It'll just take a second.

OpenStudy (anonymous):

If you look at that graph you can read off the two numbers you need.

OpenStudy (anonymous):

2, 10?

OpenStudy (anonymous):

Yup, exactly. Just the coordinates of the vertex.

OpenStudy (anonymous):

2 units, for a max profit of 10.

OpenStudy (anonymous):

But that doesn't make since These are the options for my answers: a.4,000 units; $300 b.2,000 units; $1000 c.4,000 units; $200 d.2,000 units; $200

OpenStudy (anonymous):

Oh the scale is just different. They say in the question: "weekly profit, P, in hundreds of dollars" and "number of units produced per week, in thousands" So it not 2 and 10, it's 2*1000= 2000 and 10 * $100 = $1000.

OpenStudy (anonymous):

OOOHHH! OK, I got it. Thanks. :)

OpenStudy (anonymous):

No problem.